# Supershear–subshear–supershear rupture sequence during the 2025 Mandalay Earthquake in Myanmar

### v2

@@ Abstract
We investigate the rupture dynamics of the Mw 7.7 Mandalay earthquake in Myanmar on March 28, 2025. The earthquake involved a near-vertical strike-slip rupture exceeding 400 km, with shallow slip up to 6 m. A previous inversion result revealed that the rupture initially propagated at supershear velocities ($\sim$6 km/s) near the hypocenter. A unique video recording of surface rupture, captured 124 km south of the epicenter, indicates a deceleration to subshear speeds ($\sim$3 km/s) before reaching the camera location. This deceleration is supported by observed fault-normal acceleration patterns. Satellite imagery also indicated a local minimum in slip (2–3 m) around 50 km south of the epicenter, suggesting a region of reduced stress drop, which likely caused the temporary deceleration. Beyond this point, the rupture appears to have re-established supershear propagation.
@@ 1 Introduction
Qualitatively and quantitatively describing coseismic surface displacement, deformation, and rupture in the immediate vicinity of a fault remains a critical challenge for both seismology and earthquake engineering. While strong motion records within a few hundred meters of the fault have been obtained from events such as the 2000 Western Tottori earthquake [7] and the 2023 Kahramanmaraş earthquake [15], these observational opportunities are limited due to the difficulty in predicting rupture locations in advance. Such data significantly improve the spatial resolution of a strong motion distribution and play a crucial role in understanding near-fault ground motion.
@@
At 06:20 UTC on March 28, 2025, a magnitude 7.7 earthquake occurred along the Sagaing fault, which traverses Myanmar longitudinally. The epicenter was located at $22.001^{\circ}$, $95.925^{\circ}$, with a focal depth of 10 km [21]. The event involved the rupture of a near-vertical strike-slip fault spanning approximately 400 km oriented north-south. Severe damage from strong ground motion was reported in Myanmar and Thailand, with numerous instances of surface rupture especially observed in areas along the fault trace.
@@
On May 11, 2025, a video purportedly capturing coseismic surface rupture due to the mainshock was uploaded online (https://www.facebook.com/htin.aung.33/videos/1041579804084512). The footage was taken approximately 124 km south of the epicenter. Based on the posted latitude and longitude information ($20.8821^{\circ}$, $96.0353^{\circ}$), the camera, hereafter referred to as CCTV, was located within a megawatt solar power plant adjacent to the northern part of Thazi. The shadows and timestamp within the video suggested that the camera was facing south.
@@
The video displays a timestamp (in seconds) in the upper right corner, which appears to be about five minutes behind the local standard time when compared to the USGS earthquake origin time. The video begins to subtly shake around 12:46:30 at this timestamp, with shaking intensifying from 33 seconds. Between 35 seconds and 1–2 seconds later, the footage clearly shows right-lateral strike-slip displacement of the ground beyond the fence on the right side of the screen and the gate in the foreground. This recording is likely the world’s first direct observation of surface fault rupture during an earthquake, and it provides a wealth of information from a location where no seismographs were installed.
@@
An initial analysis of the video (https://x.com/Bimaterial/status/1922157655641952512) revealed a pulse-like slip velocity function that completed within approximately two seconds. This result has since been refined through the analysis of numerous additional points within the video, indicating that the maximum slip velocity reached 3 m/s [12], or 4.5 m/s considering field observations [8].
@@
For the Myanmar earthquake, information on near-field ground motion is likely limited to the video analysis described above and the waveform data from the GEOFON station in Naypyidaw (NPW) [11]. This station is located approximately 246 km south of the epicenter and about 2.6 km west of the fault. Consequently, it is crucial to maximize the use of all available data and explore diverse avenues to understand the source behavior of this earthquake.
@@
The rupture primarily propagated southward from the hypocenter, extending over 300 km. Notably, an inversion result suggests the propagation of a supershear rupture at a speed of $\sim 6$ km/s around the hypocenter [10]. Furthermore, in the time-corrected fault-parallel displacement waveforms [11], a rapid displacement at NPW, likely associated with the passage of the rupture front, began approximately 48 seconds after rupture initiation, with no clear S-wave observed prior to this. The average velocity during this interval is $\frac{246}{48}=5.125$ km/s. Even based solely on this record, it is strongly inferred that a significant portion of the rupture propagated at supershear velocities.
@@
Conversely, at the CCTV location, which is closer to the epicenter, the rupture might propagate at subshear rather than supershear speeds. This is supported by [12], who note that the S-wave appears to arrive 1.8 seconds earlier than the rupture front at this site. While the video’s timestamp is inaccurate, which makes the exact arrival times of the S-wave and rupture at the camera location unknown, if a substantial portion of the rupture propagated subshear from the initiation point to the CCTV location, it would contradict the inversion results [10]. This suggests a potential change in rupture propagation velocity.
@@
In this paper, we present results estimating the history of surface strong ground motion and rupture propagation velocity using information from both the video and strong motion records. Specifically, insights from ground motion in the fault-normal direction will be used to demonstrate that while the fault rupture generally propagated at supershear speeds, it likely experienced a deceleration to subshear speeds just before reaching the CCTV location.
@@ 2.1 Restriction of rupture front
Herein, the origin time of the earthquake, 2025-03-28 06:20:54 (UTC), is defined as $t=0$ sec. As previously stated, the rupture arrival time at the NPW station is estimated to be $t_{N}=48$ sec. We define $r$ (km) as the epicentral distance, with the CCTV location at $r_{C}=124$ km and the NPW station at $r_{N}=246$ km. Assuming a P-wave velocity of $V_{P}=6.0$ km/s, and given that the rupture cannot propagate faster than this velocity, the rupture front at any given location where $r\leq r_{N}$ must arrive within the time window defined by $V_{P}r\leq t\leq V_{P}r+t_{N}-V_{P}r_{N}$. The lower bound of this inequality signifies that rupture cannot occur before the arrival of the P-wave radiated from the hypocenter. The upper bound accounts for the allowable delay time for the rupture to reach NPW at $t=t_{N}$.
@@
Next, we focus on the observation of a clear S-wave at the CCTV location. To investigate its duration, we analyzed the video footage, which has a resolution of 1920$\times$720 pixels and a frame rate of 30 fps. From each frame of the video, we extracted a 1-pixel-height horizontal line at 200 pixels from the top edge, spanning 1920 pixels in width. These extracted lines were then stacked vertically to construct a composite image (Fig.1).
@@
Figure 1: A composite image by extracting a 1-pixel-height horizontal slice from the video at 200 pixels from the top edge and stacking these slices vertically. The tics and values within the image correspond to the timestamp in seconds displayed in the upper right corner of the original video; the timestamp appears to be approximately five minutes behind the actual local time.
In this image analysis, it is crucial to note that the time reference relies on the seconds indicated by the timestamp in the upper right corner of the video, which differs from time $t$ as defined in this paper. We denote the seconds in the video as $T$. The precise shift between $T$ and $t$ remains unknown.
@@
Focusing on the movement of the sliding gate depicted in Fig.1, subtle shaking can be observed starting around $T=30$ sec. This is likely attributed to the P-waves originating from the hypocenter. We have to note that the exact onset cannot be determined from the image analysis and the P-wave could have arrived before $T=30$. This is because the initial phase amplitude is typically low and we cannot estimate the sensitivity of the image to ground shaking, which inherently depends on the rigidity of the structure holding the camera. Subsequently, intense shaking commences at $T\sim 33$ sec., which is interpreted as the arrival of S-waves. Following this, observations of the right side of the screen indicate that fault slip initiated around $T=35$ sec., with the slip completing within approximately two seconds.
@@
Given the earthquake origin time, the arrival times of waves and rupture at NPW, and the duration of S-wave arrival and slip at CCTV, the hypothetical rupture propagation history is estimated as depicted in Fig.2a. The solid white line represents the direct P-wave radiated at the hypocenter. The broken white line defines the condition for the rupture to reach the NPW station at $t=48$ sec.; the rupture front must always remain between these two bounds. Despite the rupture propagation velocity being nearly equal to $V_{P}$ near the rupture initiation point [10], the observation of an approximately two-second S-wave (indicated by the horizontal black bar in Fig.2a) before the rupture front arrival at the CCTV location suggests that the rupture must have decelerated to about 3.0 km/s prior to reaching CCTV. Considering these factors, it is highly probable that the rupture experienced a temporary deceleration, as depicted in Fig.2a, rather than the constant velocity propagation. While the exact duration of the P-wave at CCTV is unclear, even if the deceleration occurred closer to CCTV, it would not significantly alter the conclusion, given that the P-wave duration in the video is at least three seconds.
@@
Figure 2: (a) A potential rupture scenario illustrating supershear-subshear-supershear propagation from the epicenter to the NPW station (star, 246 km). The subshear rupture speed is 3.0 km/s. The color contour represents the slip pulse velocity with a duration of two seconds. The horizontal black bar indicates the two-second duration of the S-wave observed at the CCTV site (124 km). (b) Slip distribution along the surface fault trace derived from Sentinel-2 imagery (see Supplementary Material for details). The ordinates are common to both (a) and (b)
@@ 2.2 Fault-normal acceleration prior to the rupture
The ground motion observed at the CCTV site immediately preceding the onset of rupture suggests subshear rupture propagation. A key observation for that is the arrival of S-wave prior to the rupture [12]. The other one is the rightward displacement of the sliding gate during the approximately two-second interval between the S-wave arrival and the initiation of fault slip as seen in Fig.1. This indicates that the ground experienced substantial acceleration towards the east (left side of the screen), causing the sliding gate to be displaced apparently westward due to inertia.
Given that this is a right-lateral strike-slip fault, the rupture is expected to propagate from the north (right foreground of the screen) to the south (left background). Generally, as such slip propagates, the ground ahead of the rupture front moves towards the left of the screen for subshear rupture and towards the right of the screen for supershear rupture, as shown by 2-D steady-state pulse solutions [5] and 3-D dynamic rupture simulations [3]. However, these models do not account for abrupt rupture deceleration directly before the observation point. Therefore, to fully understand this event, it would be beneficial to interpret the results of kinematic simulations where rupture propagation velocity is controlled.
@@
We calculated the fault-normal acceleration at a distance of $r=124$ km from the epicenter with 2-D and 3-D models. The 2-D modeling was achieved by employing the integral kernel derived for 2-D velocity fields within the framework of the Boundary Integral Equation Method [19]. The 3-D modeling with a free surface and a planar strike-slip fault down to 20 km depth was done using OpenSWPC, an open-source code utilizing a finite difference method [14]. The methodology on the 2-D and 3-D calculations are detailed in the Supplementary Material. For the fault slip time history, we assumed two distinct slip pulses: uniform supershear rupture with the propagation speed of 5.125 km/s and supershear-subshear-supershear rupture sequence as illustrated in Fig.2b. Both scenarios involve a slip velocity, which is given by $\dot{D}(t)=\sin^{2}{\dfrac{\pi(t-t_{r})}{2}}H(t-t_{r})H(t_{r}+2-t)$ with a duration of two seconds, where $H(\cdot)$ is the Heaviside function. The rupture initiation time, $t_{r}$, is dependent on the epicentral distance, $r$, and follows the scenarios. While the actual duration of fault slip can exceed two seconds at various locations, with some areas peaking more than 10 seconds after rupture initiation [10], the impact of these longer durations would arrive at the CCTV site with an additional lag. Therefore, given that our current analysis focuses only on the period immediately surrounding the rupture’s arrival, these longer slip durations are disregarded.
Our modeling reveals distinct fault-normal acceleration patterns depending on the assumed rupture propagation. In the case of constant supershear rupture (Fig.3a), only a gradually increasing westward acceleration is observed prior to the onset of slip. Conversely, an abrupt change in rupture propagation velocity (Fig.3b) leads to the generation of a significant eastward acceleration commencing two seconds prior to slip initiation. This finding demonstrates robustness, having been reproduced by both 2-D and 3-D simulations. The discrepancies observed after $t\sim 27$ s are attributed to the finite fault extent in the 3-D case, where strong radiation also appears from the bottom edge of the fault. Unlike the constant velocity scenario, Fig.3b also shows discernible ground motion caused by P-waves, which is consistent with the observations from the video analysis. Therefore, these results strongly suggest that the rupture decelerated sharply just before reaching the CCTV location.

### v3

# Supershear-subshear-supershear rupture sequence during the 2025 Mandalay Earthquake in Myanmar

@@ Abstract
We investigated the rupture dynamics of the 2025 $M_{w}$7.7 Mandalay, Myanmar earthquake, using a video recording of surface rupture, strong motion recordings, waveform simulation, and satellite imagery. Our assessment, based on the S-wave observation in the video and rupture arrival time at a seismic station 246 km south of the hypocenter, suggests that rupture decelerated to subshear speeds ($\sim$3 km/s) from initial supershear propagation ($\sim$6 km/s) before reaching the camera location. This deceleration is also supported by comparison between the fault-normal acceleration patterns seen in the video and that simulated by kinematic rupture modeling. Additionally, satellite imagery indicated a local minimum in slip (2–3 m) approximately 40–60 km south of the epicenter, suggesting a region of reduced stress drop that likely caused the temporary deceleration. Beyond this point, the rupture appears to have re-established supershear propagation.
The published version is available at https://doi.org/10.26443/seismica.v4i2.1785
@@ 1 Introduction
Describing coseismic rupture, surface displacement, and deformation near a fault remains a critical challenge for both seismology and earthquake engineering. Strong motion records within a few hundred meters of the fault have been obtained from events such as the 2000 Western Tottori earthquake (Fukuyama & Mikumo,, 2007) and the 2023 Kahramanmaraş earthquake (METU-EERC,, 2023). However, such observational opportunities are limited due to the cost of installing instruments and the difficulty in predicting rupture locations, which would support deployment of stations well in advance. Therefore, it is important not to rely only on seismic recordings, but to use whatever available data to study the rupture characteristics.
@@
At 06:20 UTC on March 28, 2025, an $M_{w}$7.7 earthquake occurred along the Sagaing fault, which traverses Myanmar from north to south (Fig.S3). The epicenter was located at $22.001^{\circ}$, $95.925^{\circ}$, with a focal depth of 10 km (USGS,, 2025). The event involved the rupture of a near-vertical strike-slip fault spanning approximately 450 km(USGS,, 2025; Inoue et al.,, 2025). Severe damage was reported in Myanmar and Thailand, with numerous instances of surface rupture, especially observed in areas along the fault trace.
@@
On May 11, 2025, a video purportedly capturing coseismic surface rupture due to the mainshock was made available online (for details, see Data and code availability section, and Fig.S3 for the location). The footage was taken approximately 124 km south of the epicenter. Based on the posted latitude and longitude information ($20.8821^{\circ}$, $96.0353^{\circ}$), the camera, hereafter referred to as CCTV, was located within a megawatt solar power plant adjacent to the northern part of Thazi town.
@@
The shadows seen in the video suggested that the camera was facing southwest (Latour et al.,, 2025). The timestamps in the upper right corner of the frames appear to be about five minutes behind the local standard time when compared to the USGS earthquake origin time (USGS,, 2025). The timestamps are in hh:mm:ss format. Hereinafter, the seconds component (ss) is defined as $T$. The footage indicated that subtle shaking had begun by $T=$30, which intensified after $T=$33. Between $T=$35 and $T=$37, the footage clearly shows right-lateral strike-slip displacement of the ground beyond the fence on the right side of the video frames and the foreground gate. This recording is likely the world’s first direct observation of surface fault rupture during an earthquake. It provides direct information about the rupture at a location where no seismographs were installed.
@@
An initial analysis of the video revealed a pulse-like slip velocity function that completed within approximately two seconds (Hirano,, 2025, see Supplementary Material for the method). Although the dimension of the initial result was pixels/s, it has since been refined through the analysis of numerous additional points within the video, indicating that the maximum slip velocity reached 3 m/s (Latour et al.,, 2025; Kearse & Kaneko,, 2025), or 4.5 m/s considering field observations (Gao et al.,, 2025).
@@
For the Myanmar earthquake, information on near-field ground motion (e.g., within 10 km from the fault trace) is likely limited to the video analysis described above and waveform data from the GEOFON station in Naypyidaw (NPW) (Lai et al.,, 2025). This lone station is located approximately 246 km south of the epicenter and about 2.6 km west of the fault. Consequently, it is crucial to leverage all potential data and explore diverse avenues to understand the source behavior of this earthquake.
@@
The rupture primarily propagated southward from the hypocenter. An inversion result suggests the propagation of a supershear rupture around the hypocenter (Inoue et al.,, 2025). Furthermore, in the time-corrected fault-parallel displacement waveforms (Lai et al.,, 2025), a rapid displacement at NPW, likely associated with the passage of the rupture front, began approximately 50 s after rupture initiation, with no clear S-wave priorly observed. The average rupture velocity during this interval is $\frac{246}{50}=4.92$ km/s. Even based solely on this record, it is strongly inferred that a significant portion of the rupture propagated at supershear velocities.
@@
Conversely, at the CCTV location, which is closer to the epicenter, the rupture might propagate at subshear rather than supershear speeds. This is supported by Latour et al., (2025), who note that the S-wave appears to arrive 1.8 seconds earlier than the rupture front at this site. While the video’s timestamp is inaccurate, which makes the exact arrival times of the S-wave and rupture at the camera location unknown, if a substantial portion of the rupture propagated subshear from the initiation point to the CCTV location, it would contradict the inversion results (Inoue et al.,, 2025). Therefore, a change in the rupture propagation velocity would be consistent.
@@
In this paper, we estimate the history of ground motion and rupture propagation velocity using information from both the video and strong motion records. Specifically, we demonstrate that the fault rupture decelerated to subshear speeds just before reaching the CCTV location, although it generally propagated at supershear speed.
@@ 2.1 Restriction of rupture front
Herein, the origin time of the earthquake, 2025-03-28 06:20:52 UTC (USGS,, 2025), is defined as $t=0$ s. As previously stated, the rupture arrival time at the NPW station is estimated to be $t_{N}=50$ s. We define $r$ (km) as the epicentral distance, with the CCTV location at $r_{C}=124$ km and the NPW station at $r_{N}=246$ km. Assuming a P-wave velocity of $V_{P}=6.0$ km/s, and given that the rupture cannot propagate faster than this velocity, the rupture front at any given location where $r\leq r_{N}$ must arrive within the time window defined by $V_{P}^{-1}r\leq t\leq V_{P}^{-1}r+t_{N}-V_{P}^{-1}r_{N}$. The lower bound of this inequality signifies that rupture cannot occur before the arrival of the P-wave radiated from the hypocenter. The upper bound accounts for the allowable delay time for the rupture to reach NPW at $t=t_{N}$. If this inequality is violated at any $r(<r_{N})$, the rupture cannot arrive at $r_{N}$ at time $t=t_{N}$, even if its subsequent propagation occurs at the P-wave velocity.
@@
Next, we focus on the observation of a clear S-wave at the CCTV location. To investigate its duration, we analyzed the video footage, which has a resolution of 1280$\times$720 pixels and a frame rate of 30 fps. From each frame of the video, we extracted a 1-pixel-height horizontal line at 200 pixels from the top edge, spanning 1280 pixels in width. These extracted lines were then aligned vertically to construct a stacked image (Fig.1).
@@
Figure 1: (a) The upper part of the original video frame. The white line between the two horizontal arrows at 200 pixels from the top edge indicates the extracted line in the following process. (b) A stacked image obtained by extracting a 1-pixel-height slice along the horizontal line in (a) and aligning these slices vertically. The tics and values within (b) correspond to the timestamp in seconds displayed in the upper right corner of the original video; the timestamp appears to be approximately five minutes behind the actual local time when compared to USGS, (2025).
@@
Focusing on the movement of the sliding gate depicted in Fig.1, the subtle shaking started no later than $T=30$ s is likely attributed to the P-waves originating from the hypocenter. We have to note that the exact onset cannot be determined from the image analysis and the P-wave could have arrived before $T=30$ s. This is because the initial phase amplitude is typically low and we cannot estimate the sensitivity of the image to ground shaking, which inherently depends on the rigidity of the structure holding the camera. Subsequently, intense shaking commences at $T\sim 33$ s, which is interpreted as the arrival of S-waves. Following this, observations of the right side of the video frames indicate that fault slip initiated around $T=35$ s, with the slip completing within approximately two seconds.
@@
Given the earthquake origin time, the arrival times of waves and rupture at NPW, and the duration of S-wave arrival and slip at CCTV, the hypothetical rupture propagation history is estimated as depicted in Fig.2a. The rupture front must always remain between the two white lines in Fig.2a. Despite the rupture propagation velocity being supershear near the rupture initiation point (Inoue et al.,, 2025), the observation of an approximately two-second S-wave (indicated by the horizontal black bar in Fig.2a) before the rupture front arrival at the CCTV location suggests that the rupture must have decelerated to about 2.7–3.1 km/s at 65–99 km from the hypocenter as calculated in Supplementary Material.
@@
Figure 2: (a) A potential rupture scenario illustrating supershear-subshear-supershear propagation from the epicenter to the NPW station (triangle, hypocentral distance $r=$ 246 km). The gray-to-blue color contour represents the pulse-like slip velocity with a duration of two seconds. The solid white line represents the direct P-wave radiated from the hypocenter. The dashed white line indicates the condition under which the rupture reaches the NPW station at $t=50$ s. If the rupture front crosses to the right of this line, it cannot reach the NPW station by $t=50$ s. Solid and dotted yellow lines indicate propagation velocity of 3.5 km/s through the hypocenter and the deceleration point, respectively. The inset shows a zoomed plot where the horizontal black bar indicates the two-second duration of the S-wave before the rupture onset observed at the CCTV site ($r=$ 124 km). (b) Slip distribution along the surface fault trace derived from Sentinel-2 imagery. The ordinates are common to both (a) and (b). See Supplementary Material for algebra to obtain the deceleration point ($r_{d}=$ 65–99 km) and subshear rupture velocity (2.7–3.1 km/s) in (a) and an image analysis to calculate the slip in (b).
@@ 2.2 Fault-normal acceleration prior to the rupture
The ground motion observed at the CCTV site immediately preceding the onset of rupture suggests subshear rupture propagation. A key observation for that is the arrival of S-wave prior to the rupture (Latour et al.,, 2025). The other one is the rightward displacement of the sliding gate during the approximately two-second interval between the S-wave arrival and the initiation of fault slip as seen in Fig.1. This indicates that the ground experienced substantial acceleration towards the east (left side of the video frames), causing the sliding gate to be displaced apparently westward due to inertia.
Given that this is a right-lateral strike-slip fault, the rupture is expected to propagate from the north (right foreground of the video frames) to the south (left background). Generally, as such slip propagates, the ground ahead of the rupture front moves towards the left of the video frames for subshear rupture and towards the right of the video frames for supershear rupture, as shown by 2-D steady-state pulse solutions (Dunham & Archuleta,, 2005) and 3-D dynamic rupture simulations (Abdelmeguid et al.,, 2024). However, these models do not account for abrupt rupture deceleration directly before the observation point. Therefore, to fully understand this event, it would be beneficial to interpret the results of kinematic simulations where rupture propagation velocity matches observations from the Myanmar event.
@@
We calculated the fault-normal acceleration at a distance of $r=124$ km from the epicenter with 2-D and 3-D models. The 2-D modeling was achieved by employing the integral kernel derived for 2-D velocity fields within the framework of the Boundary Integral Equation Method (Tada & Madariaga,, 2000). The 3-D modeling with a free surface and a planar strike-slip fault down to 20 km depth was done using OpenSWPC, an open-source code utilizing a finite difference method (Maeda et al.,, 2017). The methodology on the 2-D and 3-D calculations are detailed in Supplementary Material. For the fault slip time history, we assumed two distinct slip pulses: uniform supershear rupture with the propagation speed of 4.92 km/s as calculated in the introduction and supershear-subshear-supershear rupture sequence as illustrated in Fig.2a. Both scenarios involve a slip velocity, which is given by $\dot{D}(t)=\sin^{2}{\dfrac{\pi(t-t_{r})}{2}}H(t-t_{r})H(t_{r}+2-t)$ with a duration of two seconds, where $H(\cdot)$ is the Heaviside function. The rupture initiation time, $t_{r}$, is dependent on the epicentral distance, $r$, and follows the scenarios. While the actual duration of fault slip can exceed two seconds at various locations, with some areas peaking more than 10 seconds after rupture initiation (Inoue et al.,, 2025), the impact of these longer durations would appear in the right side of the dashed white line in Fig.2a. Therefore, given that our current analysis focuses only on the period immediately surrounding the rupture’s arrival, these longer slip durations are disregarded.
Our modeling reveals distinct fault-normal acceleration patterns depending on the assumed rupture propagation. In the case of constant supershear rupture (Fig.3a), only a gradually increasing westward acceleration is observed prior to the onset of slip. Conversely, an abrupt change in rupture propagation velocity (Fig.3b) leads to the generation of a significant eastward acceleration commencing two seconds prior to slip initiation. This finding demonstrates robustness, having been reproduced by both 2-D and 3-D simulations. The discrepancies observed after $t\sim 29$ s are attributed to the finite fault extent in the 3-D case, where strong radiation also appears from the bottom edge of the fault. Unlike the constant velocity scenario, Fig.3b also shows discernible ground motion caused by P-waves, which is consistent with the observations from the video analysis. Therefore, these results suggest that the rupture decelerated prior to reaching the CCTV location.